# Next Best Sense: Guiding Vision and Touch with FisherRF for 3D Gaussian Splatting

### v3

@@ Abstract
We propose a framework for active next best view and touch selection for robotic manipulators using 3D Gaussian Splatting (3DGS). 3DGS is emerging as a useful explicit 3D scene representation for robotics, as it has the ability to represent scenes in a both photorealistic and geometrically accurate manner. However, in real-world, online robotic scenes where the number of views is limited given efficiency requirements, random view selection for 3DGS becomes impractical as views are often overlapping and redundant. We address this issue by proposing an end-to-end online training and active view selection pipeline, which enhances the performance of 3DGS in few-view robotics settings. We first elevate the performance of few-shot 3DGS with a novel semantic depth alignment method using Segment Anything Model 2 (SAM2) that we supplement with Pearson depth and surface normal loss to improve color and depth reconstruction of real-world scenes. We then extend FisherRF, a next-best-view selection method for 3DGS, to select views and touch poses based on depth uncertainty. We perform online view selection on a real robot system during live 3DGS training. We motivate our improvements to few-shot GS scenes, and extend depth-based FisherRF to them, where we demonstrate both qualitative and quantitative improvements on challenging robot scenes. For more information, please see our project page at armlabstanford.github.io/next-best-sense.
@@ I INTRODUCTION
As robots become more capable of performing manipulation in daily tasks, it is important for them to efficiently model the scene and relevant objects [1, 2]. One prominent explicit 3D representation is 3D Gaussian Splatting (3DGS) [3], which reconstructs a 3D scene from RGB images and camera poses. Traditionally, the creation of 3DGS scenes is done with a). many views and b). complete human supervision. 3DGS is well-known for requiring tens of views and near-perfect camera poses; as it is prone to overfitting on sparse input views. This approach, however, is inconsistent with many of the potential robotic applications of 3DGS. In a robotic setting, every additional view is costly, requiring movement and potentially limited onboard compute. It is often inefficient to collect hundreds of views required for training a 3DGS model in robotic environments. For future applications of 3DGS on robotics systems, robust few-view training methods are needed. Even the assistance of depth priors for regularization in prior work [4, 5, 6, 7] often relies on high quality visual data for constructing a precise 3D scene.
Figure 1: Our method outperforms random view selection for few view 3DGS scenes. We show a series of robot views, with the next view proposed by our method compared to random. Our method maximizes the Fisher information gain for each view over both vision and touch inputs.
@@
The integration of touch into 3DGS is still limited, however – only Touch-GS [14] and Snap-It, Tap-It, Splat-It [15] address touch-informed 3DGS, as compared to several works fusing vision and touch for manipulation [16, 17, 18, 19, 20, 21]. While Touch-GS demonstrates the improvements on 3DGS scenes with touch, it suffers from two shortcomings: 1. it requires a human to move the robot to desired vision and touch poses and 2. requires the order of 100-500 touches to construct a visually and geometrically sufficient scene. Snap-It, Tap-It, Splat-It similarly requires manual views. Our method automonously selects views and touches guided by uncertainty.
@@
The remainder of the paper is outlined as follows. Section II covers mathematical preliminaries, and Section III introduces our method. Section IV showcases our results, and finally, Section V discusses these results and future avenues of research.
@@ III-A1 SAM2 Depth Alignment
We posit that depth alignment should be done semantically, and supervision should be done relatively. When aligning a depth image and camera image, the depth of objects and components of a scene is based on the fact that the objects are semantically different. Concretely, we run the Segment Anything Model 2 automatic mask generator, which outputs a list of masks $M$ in image $I$. The alignment can be observed in Fig. 3. With an metric depth map from a depth sensor $D_{\text{real}}$, corresponding image $I$, monocular depth $D_{\text{MDE}}$, and list of masks $M$, we perform a mask-aware depth alignment as follows:
@@
where $s_{M_{i}}^{*}$ is the scale factor on the pixels in mask $M_{i}$, $t_{M_{i}}^{*}$ is the offset, and $p$ is a depth keypoint from sensor depth image $D_{\text{real}}$. An example of the few shot depths back-projected into 3D and into a mesh can be found in Fig. 2. Finally, the user is asked to use SAM2’s point prompt to mark the selected object in a single frame. We then propagate this to all other frames, which computes a mask. We add this mask to our list of masks to ensure object alignment. This method maintains the metric accuracy of a depth sensor but the geometric quality of monocular depth models.
@@
We lift the SAM2 aligned monocular depths into 3D and take a random percentage of them to initialize 3DGS, where a Gaussian of random color is constructed from each depth pixel. We find that this guides the few views to learn smoother geometries in the scene as compared to initialization from sensor depth data.
@@ III-A2 Geometric Depth Guidance
Drawing from prior works in few-shot GS, we introduce Pearson Relative Depth Loss to gently guide depth in our scene. We use the monocular depth output from a model such as DepthAnythingV2 [25] or Metric3DV2 [26] and measure the distribution difference between rendered and monocular depth maps.
We also incorporate scale regularization in our work, detailed in PhysGaussian [27], to reduce highly anisotropic Gaussians in our scene. Normal regularization is also lightly used, in which we use work from DN-Splatter [28] to derive normals, where we can approximate the surface normals from the monocular depth map, and during training, we minimize one of the scaling coefficients during training to encourage the Gaussians to become a flat like disc. Finally, we apply camera optimization on our views by adjusting the translation and rotation in $SO(3)\times\mathbb{R}^{3}$ for each view.
@@ III-B Next Best View for Robotics
Real world robot environments are visually feature-rich and full of color. FisherRF will select views that have a high Hessian with respect to the rendered color image, meaning that candidate views with high gradients of color are more likely to be selected. In the case of color-rich, opaque surfaces, FisherRF will select new views around these areas simply because of the high change in color. We assert that FisherRF should be extended to encode depth uncertainty, which is a strong indicator of geometric accuracy, and may be an even stronger metric of uncertainty as compared to color uncertainty. To achieve this, we formulate a depth-based negative log-likelihood objective for the radiance field similar to [13] as follows:
$-\log p(y_{D}|x,\mathbf{w})=(y_{D}-f_{D}(x,\mathbf{w}))^{T}(y_{D}-f_{D}(x,\mathbf{w}))$ (7)
where $y_{D}$ is the ground truth depth, and $f_{D}$ is the depth rasterizer that outputs depth given camera pose $x$ and Gaussian Splat parameters $\mathbf{w}$. As this objective simply replaces color terms with depth terms in FisherRF, we follow the derivation in [13] and formulate the depth information gain objective as
$\mathop{\text{argmax}}\limits_{x_{i}^{\text{acq}}}\text{tr}(\mathbf{H}^{\prime\prime}[y_{D,i}^{\text{acq}}|x_{i}^{\text{acq}},w^{*}]\mathbf{H}^{\prime\prime}[w^{*}|D^{\text{train}}]^{-1}),$ (8)
where $y_{D,i}^{\text{acq}}$ is the depth. However, the Hessian $\mathbf{H}^{\prime\prime}$ only depends on the model input and parameters, so it can be reduced to the Jacobian as $\text{diag}(\nabla_{w}f_{D}(x;w^{*})^{T}\nabla_{w}f_{D}(x;w^{*}))+\lambda I$, where we compute the gradient of the depth rasterization for a given view. We then define the information gain as a linear combination of the color and depth information gain:
$\mathcal{I}[w]=\alpha\mathcal{I}[w;f_{C}]+\beta\mathcal{I}[w;f_{D}],$ (9)
where $\alpha$ weights the color information gain, and $\beta$ weights the depth information gain, and $f_{C}$ and $f_{D}$ are the color and depth rasterizers, respectively. This addition is computationally efficient as it requires one more backward pass on depth per view. Our formulation allows to encode the uncertainty as how much depth changes in a given view; discontinuous depths in and floaters are more likely to be removed.
III-B1 Robotic Next Best View
@@ III-C Next Best Touch
We select touches similarly to vision. We compute the next best touch by computing the expected information gain of a pseudo touch camera only on depth. This ensures that more spikey and hole-like areas along an object are selected. We first only compute the depth uncertainty of Gaussians from the SAM2 object mask, ensuring spurious Gaussians not part of the object are ignored. As we add touches, we add them to the list of training views with different intrinsics to represent the touch camera. However, the touch poses are often too close to the surface of the object, which we address by moving the candidate touch view away from the object along its z axis till a surface of non-zero depth can be rendered.
We then collect a tactile image $I_{t}$ and its pose $p_{i}$. We first use a pretrained DenseNet model [29] to predict the depth along the ray, denoted as $I_{d}$. For each ray $\boldsymbol{t}_{i}$ from the fish-eye camera, we unproject pixels into the point cloud point $\boldsymbol{p}_{i}=d\cdot\boldsymbol{t}_{i}$. We filter the points by their radial distance from the center and then by their z coordinate. Directly projecting the point cloud points to camera frame will lead to a sparse depth map. We convert the point cloud to triangle mesh using the connectivity from tactile image, and then rasterize the triangles into the image plane. The depth image is later used as depth supervision for Gaussian training, where we initialize Gaussians at the point of touch. We adopt $\mathcal{L}_{1}$ depth loss with smoothing around the point of touch, and the whole process is shown in Fig. 4.
@@
TABLE II: Blender simulated few-shot ground truth depth comparison. D-ABS is the error across the whole view; D-ABS-O is the error for only the bunny.
@@
FisherRF RGB 24.41 0.68 0.22
FisherRF Depth 24.63 0.69 0.21
@@
Bunny FisherRF RGB 13.47 0.65 0.55
@@
Our Method 16.77 0.74 0.39
Bunny/Ridged Object Random 17.24 0.75 0.39
Our Method 17.47 0.76 0.38
@@ IV-A1 Stanford Bunny Blender
As seen in Table I, depth supervision guides the geometric and visual quality of 3DGS; however, it still suffers from sparsity. Lifting the aligned Gaussians improves the density of the scene but is not as geometrically accurate as using SAM2 aligned depths for initialization. Across all visual metrics, SAM2 aligned depths improve the visual quality across test scenes besides LPIPS (shown as the green value). Further, the mean absolute error in depth in the SAM2 aligned depths is the best, highlighting its usefulness as a depth prior.
@@ IV-A2 Stanford Bunny Real and Challenge Objects
With a 3D printed model of the Stanford bunny, we highlight our improvements in a real scene with low resolution visual and depth data on 8 views. In Table III, metric depth loss on aligned depths becomes infeasible, and our method outperforms current baselines, even Touch-GS [14], which uses hundreds of extra touches on the same number of views. We further include camera optimization, which improves the quality of the scene as the camera pose of the robot is not optimal. Finally, we test our method on a challenging mirror and prism on 8 views, with normal supervision and camera optimization turned on, as shown in Table IV. We refer interested readers to our website for more information.
IV-A3 Table Scene Ablation
We ablate on our few-shot Gaussian Splatting method on a challenging and colorful table scene that was constructed in real-time; namely scene 4 in Fig 6. On starting with few views, we find that the lifted SAM2 depth alignment performs the best across all metrics, and we note that it outperforms using pure depth data from the Realsense, which suffers from noise and geometric inaccuracies. Visualizations can be found on our website.
@@ IV-B Offline FisherRF
We highlight the improvement of FisherRF for depth to our two bunny scenes, reporting the mean PSNR, SSIM, and LPIPS from five trials. We train GS for 15000 steps and select a new view every 2000 steps according to FisherRF, turning camera optimization off to see the direct effect of an added view and its pose. We verify on random, FisherRF depth, RGB (baseline), and a combination of the two with our formulation in Eq. 9. Selecting random views performs the worst, and while the baseline FisherRF on color is a significant improvement over random views, it does not select views that improve the geometric construction, as seen in Fig 5, where the geometries to the right of the back chair are incorrect. In fact, the head of the bunny is the sharpest in FisherRF depth. We posit that FisherRF on depth additionally receives the benefits of FisherRF color – we notice that very different colored Gaussians, which are higher areas of color uncertainty, often are associated with ambiguous depths. However, the converse is not necessarily true, which provides additional benefit to the scene. The difference is smaller between the two in the real world; however, we choose to use FisherRF on depth for the following experiments.
@@ IV-C Robotic FisherRF
We start with only 8 random views. For each object, we keep the starting views the same for each method, and the list of candidate views every 2000 steps is kept the same (10 kinematically feasible poses). We train to 21000 steps. All experiments are run in real-time. We test each Splat on a held-out set per object of 25-40 views. The result is summarized in Table VI, with qualitative results in Fig 6. For all the objects, FisherRF performs better than the random baseline. We note that FisherRF on depth performs the best for the real world bunny – better than our baseline of FisherRF on color – which comes from FisherRF depth encouraging new views to a robot to explore the background more, which originally is highly discontinuous. We see this in the toy car as well. On the mirror, prism, and ridged objects, FisherRF performs a depth smoothing effect by suggesting new views that learn the geometry of a scene more precisely, and visually generalizes to new views. We find that real world scenes are color rich but objects are geometrically smooth. Additionally, our few-shot method is showcased, as the scene remains visually and geometrically crisp.
@@ IV-D FisherRF and Touch Data Supervision
As a final experiment, we verify the usefulness of FisherRF for single touch supervision on a mirror example. We leverage a pretrained Splat from the Touch-GS mirror and perform FisherRF depth on a list of candidate touches every 100 steps until we have added 10 touches. The result is shown in Fig 7, where more touches are selected in the hole-like areas in the mirrors, which is due to the the holes having a higher Hessian with respect to depth.
[…]
3DGS w/o Depth 16.06 0.71 0.50
Pearson Loss Depth 16.68 0.74 0.44
Lifted Depth, MSE Loss 16.30 0.73 0.45
Realsense Depth, Pearson Loss 16.83 0.73 0.47
Lifted Depth, Pearson Loss 17.05 0.75 0.38
Lifted SAM2 Depth, Pearson Loss 17.64 0.76 0.36
TABLE VII: Real world few-shot scene ablations with 3DGS on a table scene.

### v4

@@ Abstract
We propose a framework for active next best view and touch selection for robotic manipulators using 3D Gaussian Splatting (3DGS). 3DGS is emerging as a useful explicit 3D scene representation for robotics, as it has the ability to represent scenes in a both photorealistic and geometrically accurate manner. However, in real-world, online robotic scenes where the number of views is limited given efficiency requirements, random view selection for 3DGS becomes impractical as views are often overlapping and redundant. We address this issue by proposing an end-to-end online training and active view selection pipeline, which enhances the performance of 3DGS in few-view robotics settings. We first elevate the performance of few-shot 3DGS with a novel semantic depth alignment method using Segment Anything Model 2 (SAM2) that we supplement with Pearson depth and surface normal loss to improve color and depth reconstruction of real-world scenes. We then extend FisherRF, a next-best-view selection method for 3DGS, to select views and touch poses based on depth uncertainty. We perform online view selection on a real robot system during live 3DGS training. We motivate our improvements to few-shot GS scenes, and extend depth-based FisherRF to them, where we demonstrate both qualitative and quantitative improvements on challenging robot scenes. For more information, please see our project page at arm.stanford.edu/next-best-sense.
@@ I INTRODUCTION
As robots become more capable of performing manipulation in daily tasks, it is important for them to efficiently model the scene and relevant objects [1, 2]. One prominent explicit 3D representation is 3D Gaussian Splatting (3DGS) [3], which reconstructs a 3D scene from RGB images and camera poses. Traditionally, the creation of 3DGS scenes is done with a). many views and b). complete human supervision. 3DGS is well-known for requiring tens of views and near-perfect camera poses; as it is prone to overfitting on sparse input views. This approach, however, is inconsistent with many of the potential robotic applications of 3DGS. In a robotic setting, it is inefficient to collect hundreds of views, often by a human teleoperator, required for training a 3DGS model in robotic environments. For future applications of 3DGS on robotics systems, robust few-view training methods are needed. Even the assistance of depth priors in prior work on few-shot Gaussian Splatting [4, 5, 6, 7] often relies on high quality visual data for constructing a precise 3D scene, assumes that structure from motion points are already available, and require strong visual priors. Robots with noisy depth and low resolution RGB cameras cannot benefit from this.
Figure 1: Our method outperforms random view selection for few view 3DGS scenes. We show a series of robot views, with the next view proposed by our method compared to random..
@@
The integration of touch into 3DGS is still limited, however – only Touch-GS [14] and Snap-It, Tap-It, Splat-It [15] address touch-informed 3DGS, as compared to several works fusing vision and touch for manipulation [16, 17, 18, 19, 20, 21]. While Touch-GS demonstrates the improvements on 3DGS scenes with touch, it suffers from two shortcomings: 1. it requires a human to move the robot to desired vision and touch poses and 2. requires the order of 100-500 touches to construct a visually and geometrically sufficient scene. Snap-It, Tap-It, Splat-It similarly requires manual views. Our method automonously selects views and touches guided by uncertainty, in a two stage process, where vision coarsely reconstructs the object, and then touch refines it.
@@ III-A1 SAM2 Depth Alignment
We posit that depth alignment, which is done between a monocular depth image and real, noisy depth image, for Gaussian Splatting initialization, should be done semantically, and supervision should be done relatively. When aligning a real and monocular depth image and camera image, the depth of objects and components of a scene is based on the fact that the objects are semantically different. Concretely, we run the Segment Anything Model 2 automatic mask generator, which outputs a list of semantic masks $M$ in image $I$. The alignment can be observed in Fig. 3. With an metric depth map from a depth sensor $D_{\text{real}}$, corresponding image $I$, monocular depth $D_{\text{MDE}}$, and list of masks $M$, we perform a mask-aware depth alignment:
@@
where $s_{M_{i}}^{*}$ is the scale factor on the pixels in mask $M_{i}$, $t_{M_{i}}^{*}$ is the offset, and $p$ is a depth keypoint from sensor depth image $D_{\text{real}}$. An example of the few shot depths back-projected into 3D and into a mesh can be found in Fig. 2. Finally, the user is asked to use SAM2’s point prompt to mark the selected object in a single frame. We then propagate this to all other frames, which computes a mask. We add this mask to our list of masks to ensure object alignment.
@@
We lift the SAM2 aligned monocular depths into 3D and take a random percentage of them to initialize 3DGS, where a Gaussian of random color is constructed from each depth pixel. We find that this guides the few views to learn smoother geometries in the scene as compared to initialization from sensor depth data. Overall, this method maintains the metric accuracy of a depth sensor but the geometric quality of monocular depth models.
@@ III-A2 Geometric Depth Guidance
Drawing from prior works in few-shot GS [7], we introduce Pearson Relative Depth Loss to gently guide depth in our scene. We use the monocular depth output from a model such as DepthAnythingV2 [25] or Metric3DV2 [26] and measure the distribution difference between rendered and monocular depth maps.
We also incorporate scale regularization in our work, detailed in PhysGaussian [27], to reduce highly anisotropic Gaussians in our scene. Normal regularization is also lightly used, in which we use work from DN-Splatter [28] to derive normals, where we can approximate the surface normals from the monocular depth map, and during training, we minimize one of the scaling coefficients during training to encourage the Gaussians to become a flat like disc. Finally, we apply camera optimization on our views, provided in [29] by adjusting the translation and rotation in $SO(3)\times\mathbb{R}^{3}$ for each view.
@@ III-B Next Best View for Robotics
Real world robot environments are visually feature-rich and full of color. FisherRF will select views that have a high Hessian with respect to the rendered color image, meaning that candidate views with high gradients of color are more likely to be selected. We challenge this assumption that more colorful views are more uncertain. We assert that FisherRF should be extended to encode depth uncertainty, which is a strong indicator of geometric accuracy, and may be an even stronger than pure color uncertainty. To achieve this, we formulate a depth-based negative log-likelihood objective for the radiance field similar to [13], which is simply the MSE between a ground truth and rasterized depth. As this objective simply replaces color terms with depth terms in FisherRF, we follow the derivation in [13] and realize that depth information gain is nearly identical to the color information gain, but the outputs are different. We define the information gain as a linear combination of the color and depth information gain ($\mathcal{I}_{C}$ and $\mathcal{I}_{D}$):
$\mathcal{I}[w]=\alpha\mathcal{I}_{C}[w;f_{C}]+\beta\mathcal{I}_{D}[w;f_{D}],$ (7)
where $\alpha$ weights the color information gain, and $\beta$ weights the depth information gain, and $f_{C}$ and $f_{D}$ are the color and depth rasterizers, respectively. Our formulation encodes the uncertainty as the depth gradient in a given view; discontinuous depths and floaters are more likely to be removed.
@@ III-C Next Best Touch
We compute the next best touch by computing the expected information gain of a pseudo touch camera only on depth. This ensures that more spikey and hole-like areas along an object are selected. We first only compute the depth uncertainty of Gaussians from the SAM2 object mask, ensuring spurious Gaussians not part of the object are ignored. As we add touches, we add them to the list of training views with different intrinsics to represent the touch camera. However, the touch poses are often too close to the surface of the object, which we address by moving the candidate touch view away from the object along its z axis till a surface of non-zero depth can be rendered.
We then collect a tactile image $I_{t}$ and its pose $p_{i}$ from a DenseTact [30] optical tactile sensor, where a fisheye camera fires rays to a soft hemispherical silicone surface. We use a pretrained DenseNet model [31] to predict the depth along the ray from $I_{t}$, denoted as $I_{d}$. For each ray $\boldsymbol{t}_{i}$ from the tactile fish-eye camera, we unproject pixels into the point cloud point $\boldsymbol{p}_{i}=d\cdot\boldsymbol{t}_{i}$. We filter the points by their radial distance from the center and then by their z coordinate. Directly projecting the point cloud points to camera frame will lead to a sparse depth map, so we convert the point cloud to a triangle mesh using the connectivity from tactile image, and then rasterize the triangles into the image plane. The depth image is later used as depth supervision for Gaussian training, where we initialize Gaussians at the point of touch. We adopt $\mathcal{L}_{1}$ depth loss with smoothing around the point of touch, and the whole process is shown in Fig. 4.
@@
TABLE II: Blender simulated few-shot ground truth depth comparison. D-ABS (depth absolute error) is the error across the whole view; D-ABS-O is the error for only the bunny.
@@
FisherRF RGB 24.36 0.68 0.22
FisherRF Depth 24.64 0.69 0.21
@@
Bunny+Ridged Object FisherRF RGB 13.47 0.65 0.55
@@
FisherRF Depth 16.77 0.74 0.39
@@ IV-A1 Stanford Bunny Blender
As seen in Table I, depth supervision quantitatively improves the geometric and visual quality of 3DGS. We observed that lifting the aligned Gaussians improved the density of the scene, but was not as geometrically accurate as using SAM2 aligned depths for initialization. SAM2 aligned depths improve the visual quality across test scenes besides LPIPS (shown as the green value). Further, the mean absolute error in depth in the SAM2 aligned depths from Table II is the best, highlighting its usefulness as a depth prior.
@@ IV-A2 Stanford Bunny Real and Challenge Objects
With a 3D printed model of the Stanford bunny, we highlight our improvements in a real scene with low resolution visual and depth data on 8 views. In Table III, metric depth loss on aligned depths becomes infeasible, and our method outperforms current baselines, even Touch-GS [14], which uses hundreds of extra touches on the same number of views. Unlike the Blender example, we notice a significant gain in using Pearson Depth Loss as compared to simple MSE loss for depth, which we attribute to monocular depth being globally aligned with actual, noisy real world depth instead of the perturbed Blender depth.
We further include camera optimization, which improves the quality of the scene as the camera pose of the robot is not optimal. Finally, we test our method on a difficult mirror and prism on 8 views, with normal supervision and camera optimization turned on, as shown in Table IV. We still note minor gains even though these objects present a challenge for vision alone.
@@ IV-B Offline FisherRF
We highlight the improvement of FisherRF for depth to our two bunny scenes, reporting the mean PSNR, SSIM, and LPIPS from five trials in Table V. We train GS for 15000 steps and select a new view every 2000 steps according to FisherRF, turning camera optimization off to see the direct effect of an added view and its pose. We verify on random, FisherRF depth, RGB (baseline), and a combination of the two ($\alpha=0.1$, $\beta=1$) with our formulation in Eq. 7. Selecting random views performs the worst, and while the baseline FisherRF on color is a significant improvement over random views, it does not select views that improve the geometric reconstruction, as seen in Fig 5, where the geometries to the right of the back chair are incorrect. In fact, the head of the bunny is the sharpest in FisherRF depth. We posit that FisherRF on depth additionally receives the benefits of FisherRF color – we notice that very different colored Gaussians, which are higher areas of color uncertainty, often are associated with ambiguous depths. However, the converse is not necessarily true, which provides additional benefit to the scene. The difference is smaller between the two in the real world; however, we choose to use FisherRF on depth for the following experiments.
@@ IV-C Robotic FisherRF
We start with only 8 random views. For each object, we keep the starting views the same for each method, and the list of candidate views every 2000 steps is kept the same (10 kinematically feasible poses). We train to 21000 steps. We set $\alpha=0.1$, $\beta=1$ for FisherRF combined. All experiments are run in real-time. We test each Splat on a held-out set per object of 25-40 views. The result is summarized in Table VI, with qualitative results in Fig 6. For all the objects, FisherRF performs better than the random baseline. We note that FisherRF on depth performs the best for the real world bunny – better than our baseline of FisherRF on color – which comes from FisherRF depth encouraging new views to a robot to explore the background more, which originally is highly discontinuous. Additionally, FisherRF combined proves to be unusable compared to using pure depth. This suggests that color may sometimes hinder view selection, requesting to view areas that are simply colorful. Having found that FisherRF depth is the most reliable ablation, we compare it as our method to random view selection. On the car, mirror, prism, and ridged objects, FisherRF performs a depth smoothing effect by suggesting new views that learn the geometry of a scene more precisely, and visually generalizes to new views. We find that real world scenes are color rich but objects are geometrically smooth.
@@ IV-D FisherRF and Touch Data Supervision
As a final experiment, we verify the usefulness of FisherRF for single touch supervision on a mirror example. We leverage a FisherRF-depth guided Splat from the Touch-GS mirror and perform FisherRF depth on a list of candidate touches every 100 steps until we have added 10 touches. The result is shown in Fig 7, where more touches are selected in the hole-like areas in the mirrors, which is due to the the holes having a higher Hessian with respect to depth.